# CHOPS: CHat with custOmer Profile Systems for Customer Service with LLMs

### v2

@@ 5.3.1 Effectiveness and Efficiency of our proposed classifier-executor-verifier architecture
Starting from the baseline using gpt-3.5-turbo as plain Executor, we add all designed block and ablate the effectiveness of them. Finally, we substitute the Executor backbone with gpt-4 and make a comparison to plain Executor using gpt-4 as another baseline. Detailed experiment results and figure can be seen in Table 2 and Figure 5.
@@
Classifier is shown to both reduce token consumption and improve accuracy. A more complicated 2-Level Classifier can further improve accuracy while reducing cost. This free lunch can be explained as follows: By adding a classifier we can avoid sending long retrieved chunks into the Executor in some scenarios, thus improving RAG accuracy and reducing token consumption for LLMs (Chen et al., 2023). Note by shortening the chunks sent to classifier or by using weaker yet less expensive LLMs as Classifier we can save inference cost.
Self-Verification is a proven effective method to improve accuracy in previous works(Weng et al., 2023). In our experiments it is shown to improve accuracy while consuming more tokens. However, we find that in the case where gpt-4 is used for Executor, weaker and less expensive LLMs (i.e. gpt-3.5-turbo) for Verification is enough for the architecture to produce results at very good accuracy ($98\%$), while maintaining a decent cost at the same time.
In general, the proposed Classifier and Verifier are dealing with easier tasks but can effectively improve accuracy. Even in scenarios where very high accuracy ($98\%$) is required and we have to use more powerful LLMs as Executor, weaker and cheaper LLMs can still be used as Classifier and Verifier to reduce total cost while achieving satisfying accuracy.
@@ 6 Conclusion
Targeting the important scenario of Customer Service, we have collected, processed related data and proposed our CPHOS-dataset. Furthermore, we proposed CHOPS-architecture, a Classifier-Executor-Verifier agent architecture that Chat with custOmer Profile in existing Systems, offering a flexible architecture for Customer Service scenarios. Our experiments have shown that this architecture (1) improves accuracy while controlling token consumption, achieving better accuracy compared to naively using state-of-the-art LLMs, (2) provides a flexible architecture to utilize different LLMs for agent tasks with different level of requirements, thus achieving satisfying accuracy with decent cost. However, though this architecture is flexible and is not domain-specific to Customer Service in Olympiad domain and we expect it not hard to apply it to other Customer Service data, more datasets with QA pairs and Database for Customer Service is needed to further evaluate the effectiveness of our CHOPS-architecture. We hope future works may further augment our CPHOS-dataset based on the guide files, database and APIs we provide, or propose larger real-world datasets targeting the scenario of Customer Service.
Acknowledgments
We acknowledge the help on data source from CPHOS (https://cphos.cn), a non-profit entity dedicated to providing Physics Olympiad Simulations for high school students, and its participants as well. We also acknowledge members from CPHOS with whom we have had meaningful discussion, including: Xiaoyu Xiong, Xiangchen Tian, Zicheng Huang, Hongyi Liu, etc.
@@ A.1 A Real-Scene dataset: CPHOS:Cyber Physics Olympiad Simulations
In the realm of customer service, datasets such as the Customer Support on Twitter dataset Vector (2018), which gathers over 3 million tweets and replies from prominent brands on Twitter, and the Recommender Systems and Personalization Datasets McAuley , which compile a variety of user/item interactions, ratings, and timestamps, have been instrumental. The core task within customer service can be articulated as providing accurate responses or executing specific commands in response to a user’s queries or directives. This involves leveraging guiding documents or appropriately utilizing APIs. Prior datasets in file-based question answering have concentrated on reading comprehension tasks, as seen in Rajpurkar et al. (2016) and Joshi et al. (2017). Meanwhile, datasets focusing on API calling, such as Patil et al. (2023), Qin et al. (2023), and Tang et al. (2023), have primarily emphasized the use of extensive API collections for task completion, intricate reasoning, and solving mathematical problems.
@@
The CPHOS-dataset comprises a database, several guide files in PDF format, and QA pairs. These components collectively facilitate understanding how team leaders and vice team leaders can navigate the mini-program for various tasks, including uploading and grading answer sheets and accessing student grades. The QA pairs, derived from real-world interactions and further augmented by both human efforts and LLMs (including GPT-3.5 OpenAI (2022) and GPT-4 OpenAI (2023)), enrich the dataset, making it a robust resource for validation and experimentation. The dataset’s inception, modification, translation, and augmentation have been meticulously undertaken by us, with raw data sourced from CPHOS.
@@
There are 9 tables in our database, listed as following Table  3:
@@
cmf_tp_subject all answer subjects id, p_id,subject,image,grade,status,create_time
@@
MakeAllTypesToBeArbiter ChangedUserId:int change a specific user into Arbiter
GetTeacherInfoBySchoolName userId:int get user info by school name only
SchoolName:str admin user can call this successfully
ChangeAllTypesUploadLimit userId:int, Limit:int change a user’s limit for its upload limit
@@
The questions are collected from CPHOS liaison members, who are responsible for responding to (visit) team leaders. Most questions are about the use of the mini program, others are about requests to modify the marked problems or adding/modifying their status in system.
@@ A.3 Metrics
We make a comparison at the character level since different LLMs utilize different tokenization methods. Input character and output characters are separated since generating output tokens is more resource consuming than reading input tokens for LLMs (and are more expensive in terms of API price). For clarity, we define the first two metrics mathematically.
@@
A.4 More LLM Backbones
We further do experiments on the robustness of our proposed architecture one different LLMs (GLM-3, llama-2-70b). We find it hard to restrict GLM or LLaMa based verifier to output results in the regulated format, hence we set the Verifier and Classifier into gpt-3.5-turbo and mainly focuses on experimenting the effectiveness of substituting the LLM backbone for the Executor.Please refer to Table 6 for the results.

### v4

@@ 5.3.1 Effectiveness and Efficiency of our proposed classifier-executor-verifier architecture
Starting from the baseline using gpt-3.5-turbo as plain Executor, we add all designed blocks and ablate the effectiveness. Finally, we substitute the Executor backbone with gpt-4 and make a comparison to plain Executor using gpt-4 as another baseline. Detailed experiment results and figure can be seen in Table 2 and Figure 5.
@@
Classifier is shown to both reduce token consumption and improve accuracy. A more complicated 2-Level Classifier can further improve accuracy while reducing cost. This free lunch can be explained as follows. By adding a classifier we can avoid sending long retrieved chunks into the Executor in some scenarios, thus improving RAG accuracy and reducing token consumption for LLMs (Chen et al., 2023). Note by shortening the chunks sent to classifier or by using weaker yet less expensive LLMs as Classifier we can save inference cost.
Self-Verification is a proven effective method to improve accuracy in previous works (Weng et al., 2023). In our experiments, it has been shown to improve accuracy while consuming more tokens. However, we find that in the case where gpt-4 is used for Executor, weaker and less expensive LLMs (i.e. gpt-3.5-turbo) for Verification is enough for the architecture to produce results at very good accuracy ($98\%$), while maintaining a decent cost at the same time.
In general, the proposed Classifier and Verifier are dealing with easier tasks but can effectively improve accuracy. Even in scenarios where very high accuracy ($98\%$) is required and we have to use more powerful LLMs as Executor, weaker and cheaper LLMs can still be used as Classifier and Verifier to reduce total cost while achieving satisfactory accuracy.
@@ 6 Conclusion
Targeting the important scenario of Customer Service, we have collected, processed related data and proposed our CPHOS-dataset. Furthermore, we proposed CHOPS-architecture, a Classifier-Executor-Verifier agent architecture that CHat with custOmer Profile in Systems, offering a flexible architecture for Customer Service scenarios. Our experiments have shown that this architecture (1) improves accuracy while controlling token consumption, achieving better accuracy compared to naively using state-of-the-art LLMs, and (2) provides a flexible architecture to utilize different LLMs for agent tasks with different levels of requirements, thus achieving satisfying accuracy with decent cost.
However, though this architecture is flexible and is not domain-specific to Customer Service in Olympiad domain and we expect it not hard to apply it to other Customer Service data, more datasets with QA pairs and Database for Customer Service are needed to further evaluate the effectiveness of our CHOPS-architecture. We hope future works may further augment our CPHOS-dataset based on the guide files, database and APIs we provide, or propose larger real-world datasets targeting the scenario of Customer Service. We are also looking forward to future work that further improves efficiency and effectiveness in the application of Customer Service with LLMs.
Acknowledgement
The authors acknowledge the help with the data source from CPHOS (https://cphos.cn), an academic non-profit organization dedicated to providing Physics Olympiad Simulations for high school students for free. Founded in late 2020 by around 10 Physics Olympiad contestants, it now has over 100 members with 1000+ students from 100+ high schools participating in most Olympiads held. The authors also acknowledge members from CPHOS with whom they have had meaningful discussions, including: Xiaoyu Xiong, Xiangchen Tian, Zicheng Huang, Hongyi Liu, Hanyi Li, etc.
In its development process, the project acted as a course project for the NLP course advised by Professor Zhilin Yang and the TAs at Tsinghua University. The authors also acknowledge the valuable advice they provided to the project.
The authors also thank anonymous reviewers, ACs and PCs at COLM who have provided valuable suggestions that helped to improve this work.
@@ A.1 A Real-Scene dataset: CPHOS: Cyber Physics Olympiad Simulations
In the realm of customer service, datasets such as the Customer Support on Twitter dataset (Vector, 2018), which gathers over 3 million tweets and replies from prominent brands on Twitter, and the Recommender Systems and Personalization Datasets  McAuley , which compile a variety of user/item interactions, ratings, and timestamps, have been instrumental. The core task within customer service can be articulated as providing accurate responses or executing specific commands in response to a user’s queries or directives. This involves leveraging guiding documents or appropriately utilizing APIs. Prior datasets in file-based question answering have concentrated on reading comprehension tasks, as seen in Rajpurkar et al. (2016) and Joshi et al. (2017). Meanwhile, datasets focusing on API calls, such as Patil et al. (2023); Qin et al. (2023); Tang et al. (2023), have primarily emphasized the use of extensive API collections for task completion, intricate reasoning, and solving mathematical problems.
@@
The CPHOS-dataset comprises a database, several guide files in PDF format, and QA pairs. These components collectively facilitate understanding how team leaders and vice team leaders can navigate the mini-program for various tasks, including uploading and grading answer sheets and accessing student grades. QA pairs are derived from real-world interactions and are further augmented by both human efforts and LLMs (including GPT-3.5 (OpenAI, 2022) and GPT-4 (OpenAI, 2023)), which enriches the dataset, making it a robust resource for validation and experimentation. The dataset’s inception, modification, translation, and augmentation have been meticulously undertaken by us, with raw data sourced from CPHOS.
@@
There are 9 tables in our database, listed as following Table 3:
@@
cmf_tp_subject all answer subjects id, p_id, subject, image, grade, status,
create_time
@@
MakeAllTypesToBeArbiter ChangedUserId:int change a specific user
into Arbiter
GetTeacherInfoBySchoolName userId:int get user info by school
SchoolName:str name only admin user can
call this successfully
ChangeAllTypesUploadLimit userId:int, Limit:int change a user’s limit
for its upload limit
@@
The questions are collected from CPHOS liaison members, who are responsible for responding to (visit) team leaders. Most questions are about the use of the mini-program, and others are about requests to modify the marked problems or add/modify their status in system.
@@ A.3 Metrics
We make a comparison at the character level since different LLMs utilize different tokenization methods. Input characters and output characters are separated since generating output tokens is more resource-consuming than reading input tokens for LLMs (and are more expensive in terms of API price). For clarity, we define the first two metrics mathematically.
@@
A.4 Failure and Iteration Number
In experiments we repeat the iteration for at most $5$ times before giving all previous answers to a summarizer for final answer. In practice, the first iteration will give the correct answer (and also terminates) in more than half cases, two iterations significantly boost accuracy for system instructions, and three iterations are enough for file-based QAs in most cases. At Table 6 we record the number of correct (positive) and wrong (negative) answers per iteration and the transition between iterations.
iteration index 1st 2nd 3rd 4th 5th
end(TP) 151 23 9 2 4
[…]
end(FP) 3 4 3 0 7
Table 6: Number of Pos/Neg in each iteration and transition between iterations.
A.5 More LLMs
We further do experiments on the robustness of our proposed architecture on different LLMs (GLM-3, llama-2-70b). We find it hard to restrict GLM or LLaMa-based verifier to output results in the regulated format, hence we set the Verifier and Classifier into gpt-3.5-turbo and mainly focus on experimenting with the effectiveness of substituting the LLM backbone for the Executor. Please refer to Table 7 for the results.
@@ A.5 More LLMs
A specific example is LLaMA-2. During our experiments we find it hard to constrain LLaMA-2 to output in certain format: as a classifier, it tends to output multiple choices rather than one choice. As a supplement, we use the most frequent character in LLaMa output as choice and adopt LLaMa as Classifier for further experiments in Table 8.
Classifier Backbone $Acc_{sys}$ $Acc_{file}$
LLaMa-2-7b-chat 7.7 78.4
LLaMa-2-70b-chat 97.1 80.4
gpt-3.5-turbo 95.2 90.2
Table 8: LLaMa-2 as Classifier. The 70b-chat model gives relatively decent performance, but the 7b-chat model fails to correct itself given previous failure examples, resulting in a low accuracy for system instructions.